Wath is the anser to x/5-x=4

Answer by checkley71(8403)

X/5-X=4 OR X/5=4+X CROSS MULTIPLYING WE GET X=20+5X OR -4X=20 OR X=-5
PROOF -5/5-(-5)=4 OR -1+5=4 OR 4=4